# TIF-UNIMI-2025-23

### v1

@@ 8 Transverse momentum resummation
The resummation formalism that we have discussed so far is conceptually interesting and transparent, but of limited phenomenological interest, because measured cross-sections vanish at threshold, and become very small close to threshold. Hence in the region where resummation effects are significant the measured cross-section is quite small. A related resummation formalism which is instead very relevant phenomenologically is that for transverse-momentum distributions, such as the Drell-Yan cross-section, but now differential not only with respect to the gauge boson mass, but also with respect to its transverse momentum. This can be thought of as the differential counterpart of the cross-section that we discussed so far. The transverse momentum distribution if computed at fixed perturbative order diverges in the $q_{t}\to 0$ limit. This divergence is akin to the divergence of the real emission cross-section of Sect. 3.1: it is related to a soft-collinear Sudakov double-log, that can be similarly resummed to all orders.
@@
$d\Phi_{n}\propto dk_{t1}^{2}dk_{t2}^{2}\dots dk_{tn}^{2}\delta^{(2)}\left(\vec{k}_{t1}+\dots+\vec{k}_{tn}-\vec{q}_{t}\right),$ (114)
@@
$d^{2}q_{t}d^{2}k_{t1}\dots d^{2}k_{tn}\delta(\vec{k}_{t}^{1}+\dots+\vec{k}_{t}^{n}-\vec{q}_{t})$ $=d^{2}q_{t}\int\frac{d^{2}b}{(2\pi)^{2}}e^{i\vec{b}\cdot\vec{q}_{t}}d^{2}k_{t1}e^{-i\vec{b}\cdot\vec{k}_{t1}}\dots d^{2}k_{tn}e^{-i\vec{b}\cdot\vec{k}_{t}^{n}}$ (115)
@@
A second complication is related to the fact that in transverse momentum resummation the soft and collinear logs arise as logarithms of $q_{t}^{2}$ which, as we saw in Sect. 2.1, are present for all kind of parton emission, hence all partonic subchannels contribute, and not only the diagonal channel as in the case of threshold resummation. With these preliminary considerations, we are now ready to present the expression for the resummed transverse momentum distribution in the $ij$ partonic subchannel, with factorization scale set equal to $Q^{2}$:
@@
Note that the cross-section still depends on $Q^{2}$ and the dimensionless ratio $x=M^{2}/s$, and now also on $p_{t}$, and the resummation is performed in Fourier-Mellin space in order to factorize both the longitudinal and the transverse phase space, so the resummed result depends on $N$ conjugate to $x$ and $b$ conjugate to $q_{t}$.
The similarity of Eq. (8) to the threshold resummation Eq. (7.2) should be clear. The soft resummation scale is now $1/b^{2}$, and the exponent takes the form of an integral from the hard scale $Q^{2}$ to this soft scale, with a double logarithmic contribution $A^{q_{t}}$ and a single-logarithmic contribution $B^{q_{t}}$. Note that $A^{q_{t}}$ coincides with the cusp anomalous dimension only up to order $\alpha^{3}$, but beyond this order (corresponding to NNLL resummation) it differs from it. The functions $H_{ij}\left(N,\alpha_{s}(M^{2})\right)$ do not depend on the soft resummation scale $1/b^{2}$, and are thus analogous to the prefactor $C^{(c)}\left(\alpha_{s}(Q^{2})\right)$ in the threshold resummation formula. They are universal functions which do not depend on the specific process and only depend on the parton flavors $ij$.
@@ 9 Conclusion
On top of the transverse momentum resummation that we have briefly touched upon in Sect. 8, other applications of the resummation formalism, actively pursued in recent years, are related to the resummation of observables with hadronic final states and thus depend on color, specifically those that characterize jets and their substructure [Luisoni:2015xha, Marzani:2019hun]. Also, as mentioned in the introduction, in the past two decades many resummation results have been re-derived, and some derived for the first time, using the formalism of soft-collinear effective theory, in which the fundamental QCD Lagrangian is replaced by an effective Lagrangian in which the fundamental fields are decomposed as the sum of fields with different scaling properties that describe hard, soft and collinear excitations [Becher:2014oda].
@@
The Feynman diagrams included in these lectures have been produced using the FeynGame package [Harlander:2020cyh, Harlander:2024qbn, Bundgen:2025utt]
@@ Appendix A Emission from internal lines
We provide a brief summary of the classic argument from Ref. [Yennie:1961ad] that proves that in QED only radiation from external lines leads to infrared divergences.

### v2

@@ 8 Transverse momentum resummation
The resummation formalism that we have discussed so far is conceptually interesting and transparent, but of limited phenomenological interest, because measured cross-sections vanish at threshold, and become very small close to threshold. Hence in the region where resummation effects are significant the measured cross-section is quite small. A related resummation formalism which is instead very relevant phenomenologically is that for transverse-momentum distributions, such as the Drell-Yan cross-section, but now differential not only with respect to the gauge boson mass, but also with respect to its transverse momentum $q_{t}$. This can be thought of as the differential counterpart of the cross-section that we discussed so far. The transverse momentum distribution if computed at fixed perturbative order diverges in the $q_{t}\to 0$ limit. This divergence is akin to the divergence of the real emission cross-section of Sect. 3.1: it is related to a soft-collinear Sudakov double-log, that can be similarly resummed to all orders.
@@
$d\Phi_{n}\propto d^{2}k_{t1}d^{2}k_{t2}\dots d^{2}k_{tn}\delta^{(2)}\left(\vec{k}_{t1}+\dots+\vec{k}_{tn}-\vec{q}_{t}\right),$ (116)
@@
$d^{2}q_{t}d^{2}k_{t1}\dots d^{2}k_{tn}\delta(\vec{k}_{t1}+\dots+\vec{k}_{tn}-\vec{q}_{t})$ $=d^{2}q_{t}\int\frac{d^{2}b}{(2\pi)^{2}}e^{i\vec{b}\cdot\vec{q}_{t}}d^{2}k_{t1}e^{-i\vec{b}\cdot\vec{k}_{t1}}\dots d^{2}k_{tn}e^{-i\vec{b}\cdot\vec{k}_{t}^{n}}$ (117)
@@
A second complication is related to the fact that in transverse momentum resummation the soft and collinear logs arise as logarithms of $q_{t}^{2}$ which, as we saw in Sect. 2.1, are present for all kinds of parton emission, hence all partonic subchannels contribute, and not only the diagonal channel as in the case of threshold resummation. With these preliminary considerations, we are now ready to present the expression for the resummed transverse momentum distribution in the $ij$ partonic subchannel, with factorization scale set equal to $Q^{2}$:
@@
Note that the cross-section still depends on $Q^{2}$ and the dimensionless ratio $x=Q^{2}/s$, and now also on $q_{t}$, and the resummation is performed in Fourier-Mellin space in order to factorize both the longitudinal and the transverse phase space, so the resummed result depends on $N$ conjugate to $x$ and $b$ conjugate to $q_{t}$.
The similarity of Eq. (8) to the threshold resummation of Eq. (111) should be clear. The soft resummation scale is now $1/b^{2}$, and the exponent takes the form of an integral from the hard scale $Q^{2}$ to this soft scale, with a double logarithmic contribution $A^{q_{t}}$ and a single-logarithmic contribution $B^{q_{t}}$. Note that $A^{q_{t}}$ coincides with the cusp anomalous dimension only up to order $\alpha^{3}$, but beyond this order (corresponding to NNLL resummation) it differs from it. The functions $H_{ij}\left(N,\alpha_{s}(M^{2})\right)$ do not depend on the soft resummation scale $1/b^{2}$, and are thus analogous to the prefactor $C^{(c)}\left(\alpha_{s}(Q^{2})\right)$ in the threshold resummation formula. They are universal functions which do not depend on the specific process and only depend on the parton flavors $ij$.
@@ 9 Conclusion
On top of the transverse momentum resummation that we have briefly touched upon in Sect. 8, other applications of the resummation formalism, actively pursued in recent years, are related to the resummation of observables with hadronic final states and thus depend on color, specifically those that characterize jets and their substructure [12, 13]. Also, as mentioned in the introduction, in the past two decades many resummation results have been re-derived, and some derived for the first time, using the formalism of soft-collinear effective theory, in which the fundamental QCD Lagrangian is replaced by an effective Lagrangian in which the fundamental fields are decomposed as the sum of fields with different scaling properties that describe hard, soft and collinear excitations [1].
@@
The Feynman diagrams included in these lectures have been produced using the FeynGame package [10, 11, 3].
@@ Appendix A Emission from internal lines
We provide a brief summary of the classic argument from Ref. [19] that proves that in QED only radiation from external lines leads to infrared divergences.